Barkoff Enterprises, which uses the high-low method to analyze cost behavior, has determined that machine hours best explain the company's utilities cost. The company's relevant range of activity varies from a low of 600 machine hours to a high of 1,200 machine hours, with the following data being available for the first six months of the year: Month, Utilities Machine Hours

January $ 9,200 850

February 8,860 770

March 9,450 860

April 9,860 970

May 10,240 1,000

June 9,650 950

Using the high-low method, the utilities cost for Barkoff associated with 1,030 machine hours would be:
Answer #1

Variable cost per hour= (10240-8860)/(1000-770) = 6

Fixed cost = 10240-(1000*6) = 4240

The utilities cost for 1030 machine hour = (1030*6+4240) = $8360